### v1

@@ 6.3.2 Hamilton-Jacobi reachability function
The Hamilton-Jacobi (HJ) reachability analysis, a technique from robust optimal control theory, is used to guarantee constraint satisfaction in a rigorous way. HJ reachability computes a backward reachable set of a system, i.e., the set of states from which trajectories can reach some given target set (Mitchell et al., 2005; Bansal et al., 2017). For feasibility analysis, the target set is the unconstrained set, and the backward reachable set becomes the infeasible region. HJ reachability function aggregates infinite-step constraints by taking the maximum of them. Therefore, it considers the worst-case constraint violation in the entire trajectory, or the “closest distance” to the constraint boundary if no violation is going to happen. The definition of HJ reachability function is as follows.
Definition 6.6 (HJ reachability function).
@@
The constraint function $h(x)$ indicates the safety of the current state. As its extension to the whole temporal domain, $F^{\pi}(x)$ does not describe the risk level of the current state, but the worst case in the future. It is the maximizer that attempts to find the most dangerous constraint point along the whole state trajectory. Therefore, $F^{\pi}(x)\leq 0$ is equivalent to $h(x_{t})\leq 0$ for $t=0,1,\dots,\infty$.
Similar to CVF, HJ reachability function naturally satisfies a risky self-consistency condition and the optimal HJ reachability function satisfies a risky Bellman equation. The risky self-consistency condition is
$F^{\pi}(x)=\max\{h(x),F^{\pi}(x^{\prime})\}.$
The risky Bellman equation is
@@
In practical HJ reachability computation, a discount factor is needed on the right hand side of the above two equations so that they become contraction mappings, which enable the convergence of fixed point iteration (Fisac et al., 2019).
There are two ways to construct a virtual-time constraint with HJ reachability function. One way is to use a first-step constraint:
$g(x_{0|t})=F^{\pi}(x_{0|t})\leq 0,$ (9)
and $x_{i|t},i=1,\dots,n/\infty$ are unconstrained. By aggregating constraints using the maximum function, the virtual-time constraint is replaced by the first-step constraint of HJ reachability function. The other way is to use a two-step constraint:
$g(x_{i|t})=F^{k}(x_{i|t})\leq 0,i=0,1,$ (10)
and $x_{i|t},i=2,\dots,n/\infty$ are unconstrained. We omit the analysis of their corresponding feasible regions and containment relationships since it is similar to the case of CVF.
@@ 6.3.3 Constraint decay function
Constraint decay function (CDF) uses the remaining steps from the current state to constraint violation to aggregate the infinite-step constraints. Its basic idea is that if a state is infeasible, it will lead to a constraint violation in a finite number of steps. Otherwise, no constraint violation will happen in finite steps, i.e., the number of steps to constraint violation is infinite. The CDF is constructed as an exponential function with the number of steps to constraint violation as the exponent and a real number between zero and one as the base, which is formally defined as follows.
Definition 6.7 (constraint decay function).
@@
where $10<\gamma<$ 1 is the discount factor and $N$ is the number of steps to constraint violation starting from $x$ under $\pi$.
According to the property of exponential functions, as $N$ increases from zero to infinity, $F^{\pi}(x)$ decreases from one to zero. A larger value of CDF indicates that the current state is more dangerous in the sense that it is closer to constraint violation. A value of one means that the current state already violates the constraint. A value of zero means that the constraint will never be violated in the infinite horizon. In other words, $F^{\pi}(x)\leq 0$ is a necessary and sufficient condition for infinite-horizon constraint satisfaction. Therefore, CDF is a valid feasibility function of constraint aggregation type. The CDF given by Definition 6.7 is similar in form to the safety critic used in some safe RL literature, e.g., Thananjeyan et al. (Thananjeyan et al., 2021). Their difference is that they are used in different kinds of systems and for different purposes. The safety critic is used in stochastic systems for estimating the discounted probability of future constraint violation. The CDF is used in deterministic systems as a feasibility function for constructing virtual-time constraints and representing feasible regions.
Similar to other feasibility functions of constraint aggregation type, CDF naturally satisfies a risky self-consistency condition and the optimal CDF satisfies a risky Bellman equation. The risky self-consistency condition is
$F^{\pi}(x)=c(x)+(1-c(x))\gamma F^{\pi}(x^{\prime}).$
The risky Bellman equation is
@@
There are two ways to construct a virtual-time constraint with CDF. The first way is to constrain the CDF of policy to optimize in the first time step,
$g(x_{0|t})=F^{\pi}(x_{0|t})\leq 0,$ (11)
and $x_{i|t},i=1,\dots,\infty$ is unconstrained. The second way is to constraint the CDF of a fixed policy in the first two time steps,
$g(x_{i|t})=F^{k}(x_{i|t})\leq 0,i=0,1,$ (12)
and $x_{i|t},i=2,\dots,\infty$ is unconstrained. We omit the analysis of their corresponding feasible regions and containment relationships since they are similar to the case of other feasibility functions of constraint aggregation type.
7 Review of Constraint Formulations
In this section, we review several kinds of commonly used virtual-time constraint formulations. In essence, these formulations are applications of different types of feasibility functions. Therefore, we categorize them according to the feasibility functions they use, i.e., control invariant set or constraint aggregation.
@@ 7.1.1 Control barrier function
The concept of CBF originated from the field of control theory and was first proposed by Ames et al. (Ames et al., 2014) for continuous-time systems. It was later extended to discrete-time systems by Agrawal et al. (Agrawal and Sreenath, 2017). A large category of works solves OCPs with CBF constraints through online optimization, i.e., a single optimal action is computed each time the system arrives at a state. Ames et al. (Ames et al., 2016) unify CBF with control Lyapunov function (CLF) in the context of quadratic program (QP), in which performance objective is expressed by CLF and safety constraint is expressed CBF. The optimal control is obtained by solving the QP online and is ensured to keep the state in the invariant set of CBF. Nguyen et al. (Nguyen and Sreenath, 2016) propose exponential CBFs that enforce strict satisfaction of high relative degree safety constraints for nonlinear systems. They also develop a method for designing exponential CBFs based on techniques from linear control theory. Taylor et al. (Taylor et al., 2020) use a learning-based method to reduce model uncertainty in order to enhance the safety of a CBF-cerified controller. Their approach iteratively collects data and updates the controller, ultimately achieving safe behavior.
@@
Other works solve CBF-constrained OCPs in an offline manner using RL, i.e., they learn a policy that maps states to their corresponding optimal actions. Cheng et al. (Cheng et al., 2019) ensure safety of a model-free RL controller by combining it with a model-based CBF controller and online learning of unknown system dynamics. The CBF controller both guarantees safety and guides the learning process by constraining the set of explorable policies. Ohnishi et al. (Ohnishi et al., 2019) propose a barrier-certified adaptive RL algorithm, which constrains policy in the invariant set of CBF and optimizes the action-value function in this set. Their solutions to barrier-certified policy optimization are guaranteed to be globally optimal under mild conditions. Ma et al. (Ma et al., 2021) use a generalized CBF for systems with high relative degrees and mitigate the infeasibility of constrained policy optimization by an adaptive coefficient mechanism.
@@
Most of the above works handcraft CBFs as functions with known forms, while some other works synthesize CBFs using neural networks. Robey et al. (Robey et al., 2020) leverage safe trajectories generated by an expert to optimize a CBF in control affine systems. The learned CBF enjoys provable safety guarantees under Lipschitz smoothness assumptions on system dynamics. Qin et al. (Qin et al., 2020) jointly learn multi-agent control policies and CBFs in a decentralized framework. They propose a spontaneous policy refinement method to further enforce CBF conditions during testing. Yang et al. (Yang et al., 2023a) propose a safe RL algorithm that learns both policy and CBF in a model-free manner. They extend the CBF invariant loss to a multi-step version, which balances bias and variance and enhances both safety and performance of the policy.
@@ 7.1.2 Safety index
Safety index is commonly used as a safeguard for another (possibly unsafe) controller, e.g., an RL controller that solely maximizes reward performance. Zhao et al. (Zhao et al., 2021) propose a model-free RL algorithm that leverages a safety index to ensure zero constraint violation during training. This is achieved by an implicit safe set algorithm, which searches for safe control only by querying a black-box dynamic function. Ma et al. (Ma et al., 2022) simultaneously synthesize a safety index and learn a safe control policy with constrained RL. They learn the safety index by minimizing the occurrence of energy increases, which does not rely on knowledge about a prior controller.
In systems with control limits, there may be situations where it is impossible to find an action that satisfies the constraint of a safety index. Some works focus on how to synthesize a valid safety index under control limits. Wei et al. (Wei et al., 2022) propose a control-limits-aware safety index synthesis method for systems with bounded state-dependent uncertainties. They use convex semi-infinite programming to solve for a robust safe controller so that it is guaranteed to be realizable under control limits. Zhao et al. (Zhao et al., 2023c) propose a method for synthesizing safety index in general systems with control limits. They prove that ensuring the existence of safe control on a safe set boundary is equivalent to sum-of-squares programming. Zhao et al. (Zhao et al., 2023b) present an integrated dynamic model learning and safe control framework to safeguard any RL agent. They provide a design rule to construct a safety index under control limits and a probabilistic safety guarantee under stochastic dynamic models.
@@ 7.2.1 Cost value function
Plenty of existing works, mostly concentrating on the setting of CMDP, have applied CVF to handle constraints. Generally, most works follow the standard notion of CMDPs, where the constraints are imposed directly on CVF itself. Chow et al. (Chow et al., 2017) build a constraint based on the conditional value-at-risk of CVF and propose Lagrange multiplier methods for the constrained optimization problem. Ding et al. (Ding et al., 2020) also employ the primal-dual approach but update the primal variable via natural policy gradient and the dual variable via projected sub-gradient. This work is extended to an entropy-regularized case in (Ying et al., 2022). An equivalent linear formulation of the objective and the CVF-based constraint (Altman, 1999) is considered in (Bai et al., 2022) and solved by a stochastic primal-dual algorithm. To dampen the significant oscillation of state and cost value function during training when using Lagrange multiplier methods, Stooke et al. (Stooke et al., 2020) and Peng et al. (Peng et al., 2022) use PID control to update the Lagrange multiplier for a stabler intermediate performance. As et al. (As et al., 2022) use Bayesian world models to estimate an optimistic upper bound on task objective and pessimistic upper bounds on safety constraints. They use the augmented Lagrangian method to solve the constrained optimization problem based on these two bounds.
@@
Besides Lagrangian-based methods, other approaches have also been proposed for solving OPCs with CVF constraints. Liu et al. (Liu et al., 2022) deal with instability issues of primal-dual style methods by introducing the Expectation-Maximization approach. By adopting a non-parametric variational distribution, the constrained optimization problem in the expectation step becomes convex and can be solved analytically. In (Liu et al., 2020), the authors introduce the interior-point method to augment the objective with logarithmic barrier functions composed with CVF. Chow et al. (Chow et al., 2018) propose a Lyapunov-based approach for transient MDPs which constructs Lyapunov functions w.r.t an undiscounted version of CVF. Achiam et al. (Achiam et al., 2017) construct constrained optimization problems on the basis of a novel bound on the difference in cumulative rewards or costs between two policies and solve them with trust region methods. However, these optimization problems may be infeasible, which undermines the theoretical monotonicity. Yang et al. (Yang et al., 2020) address this by first applying an unconstrained trust region method and then projecting the policy back onto the constrained set. Likewise, Zhang et al. (Zhang et al., 2020) also propose a two-stage algorithm that first searches for a solution of a constrained optimization problem in the non-parameterized policy space and then projects it back into the parametric one. Yu et al. (Yu et al., 2022b) propose a two-policy method where a safety editor, serving as an extension of a safety shield, is trained to reconcile possible constraint violation with minimal influence on the objective. Zhao et al. (Zhao et al., 2023a) introduces the framework of maximum MDP, which is an extension of CMDP that constrains the expected maximum state-wise cost along a trajectory. Under this framework, they propose state-wise constrained policy optimization (SCPO) algorithm, which provides guarantees for state-wise constraint satisfaction in expectation.

### v2

@@ 6.3.2 Hamilton-Jacobi reachability function
The Hamilton-Jacobi (HJ) reachability analysis, a technique from robust optimal control theory, is used to guarantee constraint satisfaction in a rigorous way. HJ reachability computes a backward reachable set of a system, i.e., the set of states from which trajectories can reach some given target set [mitchell2005time, bansal2017hamilton]. For feasibility analysis, the target set is chosen as the complement of the constrained set, and the backward reachable set becomes the infeasible region. HJ reachability function aggregates infinite-step constraints by taking the maximum of them. In other words, it considers the worst-case constraint violation in the entire trajectory or the “closest distance” to the constraint boundary if no violation is going to happen. The definition of HJ reachability function is as follows.
Definition 6.3.2 (HJ reachability function).
@@
It follows directly from the above definition that $F^{\pi}(x)\leq 0$ is equivalent to $h(x_{t})\leq 0$ for $t=0,1,\dots,\infty$. Similar to CVF, HJ reachability function naturally satisfies a risky self-consistency condition:
$F^{\pi}(x)=\max\{h(x),F^{\pi}(x^{\prime})\},$
and the optimal HJ reachability function satisfies a risky Bellman equation:
@@
In practical HJ reachability computation, a discount factor is introduced to the right-hand side of the above two equations to obtain contraction mappings for fixed point iteration [fisac2019bridging].
Similar to CVF, there are two ways to construct a virtual-time constraint with HJ reachability function. One is to use a first-step constraint:
$g(x_{0|t})=F^{\pi}(x_{0|t})\leq 0,$ (6.5)
and the other is to use a two-step constraint:
$g(x_{i|t})=F^{k}(x_{i|t})\leq 0,i=0,1,$ (6.6)
We omit the analysis of their corresponding feasible regions and containment relationships since it is similar to the case of CVF.
@@ 6.3.3 Constraint decay function
Constraint decay function (CDF) uses the remaining steps from the current state to the first constraint violation to aggregate constraints [yang2024synthesizing]. Its basic idea is that if a state is infeasible, it will lead to a constraint violation in finite steps. CDF is formally defined as follows.
Definition 6.3.3 (Constraint decay function).
@@
where $0<\gamma<1$ is the discount factor, and $N$ is the number of steps to the first constraint violation starting from $x$ under $\pi$.
As $N$ increases from zero to infinity, $F^{\pi}(x)$ decreases from one to zero. A larger value of CDF indicates that the current state is more dangerous in the sense that it is closer to constraint violation. A value of one means that the current state already violates the constraint, and a value of zero means that the constraint will never be violated. In other words, $F^{\pi}(x)\leq 0$ is a necessary and sufficient condition for infinite-horizon constraint satisfaction. Therefore, CDF is a valid CA-based feasibility function. CDF is similar to a notion called safety critic in safe RL literature [thananjeyan2021recovery]. Their difference is that they are used in different kinds of systems and for different purposes. The safety critic is used in stochastic systems for estimating the probability of future constraint violation. In contrast, CDF is used in deterministic systems as a feasibility function for constructing virtual-time constraints and representing feasible regions.
Similar to other CA-based feasibility functions, CDF naturally satisfies a risky self-consistency condition:
$F^{\pi}(x)=c(x)+(1-c(x))\gamma F^{\pi}(x^{\prime}),$
and the optimal CDF satisfies a risky Bellman equation:
@@
There are also two ways to construct virtual-time constraints with CDF. The first is a single-step one:
$g(x_{0|t})=F^{\pi}(x_{0|t})\leq 0,$ (6.7)
and the second is a two-step one:
$g(x_{i|t})=F^{k}(x_{i|t})\leq 0,i=0,1.$ (6.8)
Their corresponding feasible regions and containment relationships are similar to other CA-based feasibility functions and are thus omitted here.
Chapter 7 Review of Constraint Formulations
In this chapter, we review several commonly used virtual-time constraint formulations. In essence, they are applications of different feasibility functions. Existing surveys have extensively covered specific types in isolation, such as the theory and applications of CBFs [ames2019control] or HJ reachability analysis [bansal2017hamilton, ganai2024hamilton]. Meanwhile, other reviews on safe RL [garcia2015comprehensive, gu2024review] and safe control [brunke2022safe] discuss safety constraints from a broader, algorithmic perspective, often without dissecting the unified mathematical essence of the underlying feasibility functions. In contrast, we provide a synthesizing framework that moves beyond these isolated or high-level discussions. We categorize the constraint formulations according to their core feasibility function (CIS or CA), demonstrating how they all conform to the same general mathematical structure and feasibility analysis presented in our unified framework.
@@ 7.1.1 Control barrier function
The concept of CBF originated from the field of control theory and was first proposed by ames2014control for continuous-time systems. It was later extended to discrete-time systems by agrawal2017discrete. A large category of works solves OCPs with CBF constraints through online optimization, i.e., a single optimal action is computed each time the system arrives at a state. These works formulate the virtual-time OCP online at each step, taking the current state as the initial state, and constraining the action to ensure subsequent states remain in the invariant set of the CBF. ames2016control unify CBF with control Lyapunov function (CLF) in the context of quadratic program (QP), in which performance objective is expressed by CLF and safety constraint is expressed CBF. The optimal control is obtained by solving the QP online and is ensured to keep the state in the invariant set of CBF. nguyen2016exponential propose exponential CBFs that enforce strict satisfaction of high relative degree safety constraints for nonlinear systems. They also develop a method for designing exponential CBFs based on techniques from linear control theory. Another method to deal with high relative degree is HOCBF [xiao2019control, xiao2021high], which takes high-order time derivatives of the constraint function so that where the action explicitly appears in the CBF. To address the over-conservativeness of CBFs, xiao2023barriernet replace the set of hard constraints HOCBFs with a set of differentiable constraints. This is achieved by a barrier layer that can be combined with any neural network controller. taylor2020learning use a learning-based method to reduce model uncertainty in order to enhance the safety of a CBF-cerified controller. Their approach iteratively collects data and updates the controller, ultimately achieving safe behavior.
@@
Other works solve CBF-constrained OCPs in an offline manner using RL, i.e., they learn a policy that maps states to their corresponding optimal actions. These works integrate CBF constraints into the policy optimization of the RL training process, ensuring that the learned policy renders the zero-sublevel set of the CBF forward invariant. cheng2019end ensure safety of a model-free RL controller by combining it with a model-based CBF controller and online learning of unknown system dynamics. The CBF controller both guarantees safety and guides the learning process by constraining the set of explorable policies. ohnishi2019barrier propose a barrier-certified adaptive RL algorithm, which constrains policy in the invariant set of CBF and optimizes the action-value function in this set. Their solutions to barrier-certified policy optimization are guaranteed to be globally optimal under mild conditions. ma2021model use a generalized CBF for systems with high relative degrees and mitigate the infeasibility of constrained policy optimization by an adaptive coefficient mechanism. marvi2021safe augment the cost function with a CBF and solve the resulting problem using a model-free RL algorithm, yielding a controller with the ability of proactive safety planning. emam2022safe frame safety as a differentiable robust CBF layer in a model-based RL framework, within which the underlying reward-driven task is modularly learned independent of safety constraints.
@@
Most of the above works handcraft CBFs as functions with known forms, while some other works synthesize CBFs using neural networks. robey2020learning leverage safe trajectories generated by an expert to optimize a CBF in control affine systems. The learned CBF enjoys provable safety guarantees under Lipschitz smoothness assumptions on system dynamics. qin2020learning jointly learn multi-agent control policies and CBFs in a decentralized framework. They propose a spontaneous policy refinement method to further enforce CBF conditions during testing. yang2023model propose a safe RL algorithm that learns both policy and CBF in a model-free manner. They extend the CBF invariant loss to a multi-step version, which balances bias and variance and enhances both safety and performance of the policy. To handle control input limits, liu2023safe propose a learner-critic architecture, where the critic finds counterexamples of input saturation and the learner optimizes the neural CBF to eliminate them. To address the conservativeness of learning-based CBFs, yang2024synthesizing propose feasible region iteration that learns the maximum feasible region with a CDF to generate accurate feasibility labels for learning the CBF.
@@ 7.1.2 Safety index
SI is commonly used as a safeguard for another (possibly unsafe) controller, e.g., an RL controller that solely maximizes reward performance. zhao2021model propose a model-free RL algorithm that leverages an SI to ensure zero constraint violation during training. This is achieved by an implicit safe set algorithm, which searches for safe control only by querying a black-box dynamic function. ma2022joint simultaneously synthesize an SI and learn a safe control policy with constrained RL. They learn the SI by minimizing the occurrence of energy increases, which does not rely on knowledge about a prior controller. To deal with systems with varying dynamics, yun2025safe propose an SI adaptation method that guarantees forward invariance and finite convergence of the safe control law. chen2024real leverage determinant gradient ascend and derive a closed-form update to SI parameters once the dynamics are perturbed. wei2022safe propose a method for safety control synthesis with neural network dynamic models. The method first synthesizes an SI offline using evolutionary methods, and then solves a constrained optimization problem online based on an encoding of neural networks.
In systems with control limits, there may be situations where it is impossible to find an action that satisfies the constraint of an SI. Some works focus on how to synthesize a valid SI under control limits. wei2022persistently propose a control-limits-aware SI synthesis method for systems with bounded state-dependent uncertainties. They use convex semi-infinite programming to solve for a robust safe controller so that it is guaranteed to be realizable under control limits. zhao2023safety propose a method for synthesizing SI in general systems with control limits. They prove that ensuring the existence of safe control on a safe set boundary is equivalent to sum-of-squares programming. zhao2023probabilistic present an integrated dynamic model learning and safe control framework to safeguard any RL agent. They provide a design rule to construct an SI under control limits and a probabilistic safety guarantee under stochastic dynamic models. chen2024safety study the problem of synthesizing feasible safety indices under state-dependent control spaces, and leverage Positivstellensatz to formulate the problem as nonlinear programming.
@@ 7.2.1 Cost value function
Plenty of existing works, mostly concentrating on the setting of CMDP, have applied CVF to handle constraints. Generally, most works follow the standard notion of CMDPs, where the constraints are imposed directly on CVF itself. These works typically consider the cumulative constraint, which requires the sum of costs to be less than a given threshold. When the threshold is set to zero, the cumulative constraint is equivalent to requring all costs to be zero, and becomes the statewise constraint considered in this paper. chow2017risk build a constraint based on the conditional value-at-risk of CVF and propose Lagrange multiplier methods for the constrained optimization problem. ding2020natural also employ the primal-dual approach but update the primal variable via natural policy gradient and the dual variable via projected sub-gradient. This work is extended to an entropy-regularized case in [ying2022dual]. An equivalent linear formulation of the objective and the CVF-based constraint [altman1999constrained] is considered in [bai2022achieving] and solved by a stochastic primal-dual algorithm. To dampen the significant oscillation of state and cost value function during training when using Lagrange multiplier methods, stooke2020responsive and peng2022model use PID control to update the Lagrange multiplier for a stabler intermediate performance. as2022constrained use Bayesian world models to estimate an optimistic upper bound on task objective and pessimistic upper bounds on safety constraints. They use the augmented Lagrangian method to solve the constrained optimization problem based on these two bounds.
@@
Besides Lagrangian-based methods, other approaches have also been proposed for solving OPCs with CVF constraints. liu2022constrained deal with instability issues of primal-dual style methods by introducing the Expectation-Maximization approach. By adopting a non-parametric variational distribution, the constrained optimization problem in the expectation step becomes convex and can be solved analytically. liu2020ipo introduce the interior-point method to augment the objective with logarithmic barrier functions composed with CVF. chow2018lyapunovbased propose a Lyapunov-based approach for transient MDPs which constructs Lyapunov functions w.r.t an undiscounted version of CVF. achiam2017constrained construct constrained optimization problems on the basis of a novel bound on the difference in cumulative rewards or costs between two policies and solve them with trust region methods. However, these optimization problems may be infeasible, which undermines the theoretical monotonicity. yang2020projection address this by first applying an unconstrained trust region method and then projecting the policy back onto the constrained set. Likewise, zhang2020first also propose a two-stage algorithm that first searches for a solution of a constrained optimization problem in the non-parameterized policy space and then projects it back into the parametric one. yu2022towards propose a two-policy method where a safety editor, serving as an extension of a safety shield, is trained to reconcile possible constraint violation with minimal influence on the objective. liu2023constrained study the offline safe RL problem and propose the constrained decision transformer. They train the policy in an autoregressive manner, making it adaptable to different thresholds of the cumulative cost constraints during deployment. zhang2024cvar consider the constraint of the conditional value-at-risk (CVaR) of cumulative costs, and propose the CVaR-CPO algorithm, which learns a distributional CVF to provide a quantile-based estimation of the CVaR constraint.
While the CVF is originally proposed for cumulative constraints in CMDP, the connection between CVF and state-wise constraints has been revealed by existing works. sootla2022saute propose a framework that converts the cumulative costs constraint into state-wise constraints by augmenting the state with the remaining safety budget, preserving the Markov property. zhao2023state introduces the framework of maximum MDP, which is an extension of CMDP that constrains the expected maximum state-wise cost along a trajectory. Under this framework, they propose state-wise constrained policy optimization algorithm, which provides guarantees for state-wise constraint satisfaction in expectation.